OpenStudy (anonymous):

Solving Equations by multiplying or dividing 6. 64= -4n 7. -114= 12q 8. r/11 = -12 9. -5g= -385 10. -16x= -176 P.S. the "/" means its a fraction HELP ME PLZ

OpenStudy (shamim):

4n=-64 Right?

Mehek (mehek14):

6) divide \(\dfrac{64}{-4}\)

OpenStudy (anonymous):

@Godlovesme

OpenStudy (anonymous):

yes

Mehek (mehek14):

7) \(\dfrac{-114}{12}\)

Mehek (mehek14):

8) \(-12*11\)

Mehek (mehek14):

9) \(\dfrac{-385}{-5}\)

OpenStudy (anonymous):

so if it has a fraction in it i multiply?

Mehek (mehek14):

10) \(\dfrac{-176}{-16}\)

Mehek (mehek14):

if it's division, to solve you multiply if it's multiplication, you solve by division

OpenStudy (anonymous):

ok

Mehek (mehek14):

all the questions except for 8 are being multiplied so you divide for them 8 is being divided so multiply to solve for it